4) My great-uncle died. In his will, he asked me so sell his house in Beverly Hills and share the proceeds with his nephew and with his favorite charity. Of course I care about myself but I also care about his nephew and about charity if I receive M and the nephew receives N and charity receives C, I will derive utility M12N1/3C1/6. I sell the house for $1.2 Million (a) How much money should I keep for myself, how much should I give to the nephew and how much should I give to charity?

(b) I forgot that I have to pay income tax of 50% on all the money I receive and the nephew has to pay income tax of 40% on all the money he receives. (I live in California and he lives in Nevada, so I have to pay state income tax and he does not.) After taking taxes into account, how much money should I keep for myself, how much should I give to the nephew and how much should I give to charity? (Hint: What is the price of money I receive? What is the price of money the nephew receives? What is the price of money the charity receives?)

Hint: What is the price to me of money my nephew receives?

Textbook: Nicholson & Snyder, Microeconomic Theory, 12th edition.
Answer #1

House sell 1200000$

Ratio in which 1200000$ div

ide 1/2,1/3,1/6

3:2:1is ratio in which amount divide

So M =3/6*1200000$=600000$

N=2/6*1200000$=400000$

C=1/6*1200000$=200000$

  • B)M recieve 600000$ from house sell and give 50 %tax

After tax he receive 50 /100*600000=300000

  • N receive 400000$from house sell and give 40 %tax

40 /100*400000=1600000

After tax he receive 1600000

  • Charityreceive =200000